# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ Abstract
As shown in Fig. 4(d,h), in the presence the two-particle gain $\kappa$, the onsite two-particle excitation is strongly bound during the dynamical evolution for the double-occupied state initially prepared at the odd site. For small |$\absolutevalue{\kappa}$ | with $\kappa=-0.5$, as expected, the doublon transports towards the right boundary [see Fig. 4(d)]. While, for the large two-particle gain with $\kappa=-2$, the evolution direction is reversed [see Fig. 4(h)], indicating the reversal of NHSE controlled by the two-particle gain.
Conclusion.—We investigate the loss or gain controlled NHSEs, with the combination of the Bose-Hubbard interaction, in a 1D zigzag lattice subject to a magnetic flux threading through each triangular plaquette. By introducing imbalanced onsite single-particle loss on the odd and even sites, we realize all the states, from two doublon bands, localized at the same boundary due to the interplay of the magnetic-flux-induced chiral transport and staggered single-particle loss. Imposing the uniform onsite two-particle loss or gain produces an unexpected reversal of the localization direction of one doublon band due to the effectively strong nonreciprocal hopping contributing from the virtual second-order and third-order hopping processes of particle pairs. In experiments, our predictions could be tested experimentally using ultracold atoms Liang et al. (2022); Ren et al. (2022); Cai et al. (2019) and dissipative photonic resonator Leefmans et al. (2024), and the dynamical skin-mode localization and its reversal can be observed via introducing incoherent two-particle pumping. There can also be simulated in the electrical circuits by mapping the 1D two-particle excitation into a 2D lattice Olekhno et al. (2020). In the future, it is interesting to explore the NHSEs controlled by many-body interaction and loss beyond double-excitation subspace.

### v2

@@ Abstract
As shown in Fig. 4(d,h), with the two-particle gain $\kappa$, the onsite two-particle excitation is strongly bound during the dynamical evolution for the double-occupied state initially prepared at the odd site. For small |$\absolutevalue{\kappa}$ | with $\kappa=-0.5$, as expected, the doublon transports towards the right boundary [see Fig. 4(d)]. While, for the large two-particle gain with $\kappa=-2$, the evolution direction is reversed [see Fig. 4(h)], indicating the reversal of NHSE controlled by the two-particle gain.
Conclusion.—We investigate the loss or gain controlled NHSEs, with the combination of the Bose-Hubbard interaction, in a 1D zigzag lattice subject to a magnetic flux threading through each triangular plaquette. By introducing imbalanced onsite single-particle loss on the odd and even sites, we realize all the states, from two doublon bands, localized at the same boundary due to the interplay of the magnetic-flux-induced chiral transport and staggered single-particle loss. Imposing the uniform onsite two-particle loss or gain produces an unexpected reversal of the localization direction of one doublon band. This results from the effectively strong nonreciprocal hopping contributing from the virtual second-order and third-order hopping processes of particle pairs. In experiments, our predictions could be tested experimentally using ultracold atoms Liang et al. (2022); Ren et al. (2022); Cai et al. (2019), and simulated in the electrical circuits by mapping the 1D two-particle excitation into a 2D lattice Olekhno et al. (2020). In the future, it is interesting to explore the NHSEs controlled by interaction and loss beyond double-excitation subspace.